OpenStudy (anonymous):

A rectangular plot of land is to be fenced in using two kinds of fencing. Two opposite sides will use heavy-duty fencing selling for $5 a foot, while the remaining two sides will use standard fencing selling for $3 a foot. What are the dimensions of the rectangular plot that will minimize the cost of fencing to enclose 6000 square feet?

OpenStudy (anonymous):

Area: x y = 6,000 Perimeter: P = 2 ( x + y ) Cost: C = 2 ( $3x + $5 y ) = 2 ( 3x + 5 * 6,000/ x) -> C' = .... Minimize the cost happen when C' = 0 ==> x dimension --> y

OpenStudy (anonymous):

so C' = 2(3-30000x^-2) x=100? y=60?

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

ok thanks. that just helped me out a lot. i was confused b/c i had it equal to dC/dX